UT Sep 07 2022 - Night #6 of Restart




Night Info

  • Sunset 18:38
  • Nautical Twilight Ends 19:32
  • Astro Twilight Ends 20:02
  • Astro Twilight Begins 04:34
  • Nautical Twilight Begins 05:04
  • Sunrise 05:58
  • Moonset 02:40
  • Moon illumination ~85%

Staff

  • OSA: J.Williams
  • ISp: J.Schuch
  • Mtn Manager: T.Castro
  • ISA/SO: J.Power, B.Rothberg, R.T.Gatto
  • Technical Expert: J.Hill
  • AO: D. Miller

Night Plan

ALTA prediction for Tonight

20220906_see_evol_time_Height_20_20000.png

Starting Config LBCs

LBC Skyflats in Twilight

Telescope Wakeup

  • Reconfigure for Bino-LUCIs when done with skyflats
  • Conifrmatory Field Aberration (2 hrs) - Critical (1hour) (SX more important than DX due to SX removal)
    • Requires preliminary analysis to be completed
    • Can be done in parallel (data collection, analysis in serial)
    • need both AGW1 and AGW2 on the telescope and checked to execute
    • must complete SH spot centering and hotspot check before executing
    • 1.2" or better seeing (1" or better ideal)
    • Target Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0

LUCI Wakeup

  • Pending LUCI issues:
    • LUCI-1 no MOS masks
    • Some flexure (spectroscopy) stll pending
  • LUCI Scripts Location: Service_Observing/Restart_2022/LUCI
    • LUCI_2022_Restart_V1.80.xml - all scripts (except LS Rot ZP)
    • RotZP _v2.xml - for LS RotZP scripts
    • LUCI_Phot_ZP_OT_V6.xm - LUCI photometric standards
  • rotator zeropoint (requires MOS mask) - (0.5 hr)
    • 1.2" or better seeing needed
    • LUCI-1 is empty - use the LS method
    • LUCI-2 - tentatively OK - the last two masks are consistent with < 0.1, but the last mask was done in poor seeing Take again with LS to check
    • Masks:
      • ID 903033 - Stone O
      • ID 921934 - Stone A
      • iD 931027 - Stone D
      • Do the LS method Bino to compare L2 with mask and LS method
    • Do a MOS acquisition (imaging of mask and thru-mask image confimration) - verify the theta < 0.1 deg for each mask
    • To update results (if needed)
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • Longslit Method (backup in case mask exchange(s) aren't completed in time
      • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
      • use pseudo-LMS files below for acquisition - works in similar manner to MOS acquisition - will need to manually load the LMS files in RTD
  • Collect transform data - (2hrs) - must do RotZP First
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
  • If seeing is stable and 1.2" or better - skip to AO/ESM - otherwise continue
  • Imaging check (1hr)
    • ANY seeing OK
    • dither in-field
      • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am
    • dither out-of-field (run one)
      • M52 - JHK, dither 5' out of field - visible all night
      • M40 - Broad filters, few narrow filters - visible all night
  • G210 N1.8 LS Testing (Flexure and wavelength are completed)
    • Choose one
      • NGC 6052 - twilight-22hr
      • Arp 230 23:30-twilight
  • LUCI Photometric Standards
    • Any conditions - must be clear
    • Pick any standard thats visbile, please try to get 2 diferent airmasses

AO/ESM Checkout On-Sky

  • Full AO
    • Scripts: /home/lbto/Service_Observing/Restart_2022/LUCI/AO/AO_Testing_V94.xml
    • OSCO
    • Bin 1
    • Bin 2
    • Bin 3
    • Bin 4
  • ESM
    • Bin 1
    • Bin 2
    • Bin 4

LBC Checkout Continued

  • Superfocus Sequence
    • Requires 1.2" or better stable seeing

End of Night finish w/LBCs (if clear) - Twilight Flats

Googly Doc via J.Power

Night Log

Closed Dome

  • UT 00:39 - test preset, setting up instruments (LUCI, LBCs) for night
LUCI Software problem - major issue - FlexComp will not engage - and freezes instrument manager - makes LUCIs not usable for science and impacts restart

-- BarryRothberg - 06 Sep 2022
Topic revision: r5 - 07 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback